One card is selected at random from a standard deck, not replaced, and then a second card is drawn. Find the probability of selecting two face cards
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

One card is selected at random from a standard deck, not replaced, and then a second card is drawn. Find the probability of selecting two face cards
----------------
There are 12 face cards.
---
P(1st face card) = 12/52
---
P(2nd face card) = 11/52
-----
P(2 face cards selected without replacement) = (12*11/52*51)
---
That is the same as 12C2/52C2
==================================
Cheers,
Stan H.

Answer by Alan3354(69443)

There are 12 face cards, 4 each King, Queen and Jack.
---
The 1st is 12/52
The 2nd is 11/51
--> (12*11/52*51)
= 11/221